OpenStudy (anonymous):

3 vertices of a parallelogram: (2,7)(6,1) and (9,6) Find the fourth vertex (USE ANALYTIC METHOD) Helpppp! :D

OpenStudy (anonymous):

By plotting, you might find that since it is a parallelogram, then:  a quick method is that since for B and C, compared to A and D, 9-6=x-2(their x distance must be same.) 6-5=y-7(their y distance must be same.

OpenStudy (anonymous):

Oh my. Thank you so much!! :D

OpenStudy (anonymous):

lol You're welcome :)